# What is Fair? Defining Fairness in Machine Learning for Health

### v3

@@ Abstract
There are myriad potential applications of machine learning (ML) for health, including automated disease detection, computer-aided diagnosis, and personalized treatment planning [1]. However, there is substantial evidence that, without appropriate forethought and planning, ML models can introduce or exacerbate health inequities by making less accurate decisions for certain groups or individuals [2]. Within medical imaging, state-of-the-art ML models used for disease diagnosis, risk prediction, and triage management are known to underperform within minority groups defined by protected attributes, including sex, race, and ethnicity [3, 6, 7, 8, 9, 10, 11, 12, 13, 4, 5]. For example, deep learning models used to detect 14 common diseases from chest X-rays substantially were found to under-diagnose intersectional under-served subgroups, such as Hispanic female patients, potentially resulting in treatment delays if deployed in practice [11]. Similarly, ML models trained on electronic health record (EHR) and claims data have demonstrated biased performance by underperforming in specific subgroups [14, 15]. For example, in a landmark study, a widely used commercial risk prediction model for identifying patients with complex health needs was found to have significantly lower accuracy for Black patients compared to White patients, making Black patients less likely to be recommended for essential care services than equally healthy white patients [16].
@@
As applications of ML in health become commonplace, it is crucial to recognize, account for, and mitigate such disparities in model performance in order to ensure health equity. Broadly, an ML model is said to be fair if it does not discriminate against an individual or group [17]. Concepts of fairness have been extensively studied across various disciplines, including social choice theory, game theory, economics, and law [18, 19, 20, 21, 22, 23, 24]. Building on these principles, the subfield of fairness within ML provides a framework for evaluating and mitigating bias throughout the model development process.
@@
Although fairness has been an extremely active area of research over the past decade, evaluating and implementing fairness within ML for health is relatively nascent [25, 26, 27]. A systematic review of articles utilizing ML for EHR-based phenotyping found that only 5% of studies assessed fairness [28]. Similarly, a review of EHR-based prediction models found that most studies investigate the overall performance of ML models, but do not interrogate potential biases [29]. Beyond EHR applications, several scoping reviews of clinical ML models found that the adoption of fairness remains inconsistent, partly due to knowledge gap between ML and clinical researchers [27, 30, 31]. Moreover, over two dozen definitions of fairness have been proposed, most of which originate outside health-focused literature, making it a particularly challenging domain to navigate.
@@
To review this fundamental topic, we examine key notions of fairness and their use in ML in health [^1] This review focuses on quantitative measures of fairness. Alternative approaches are discussed in the outlook section.. We first introduce biases that can emerge throughout the model development process using examples from the literature to illustrate why ML models can be unfair. We then review what it means for a model to be fair, beginning with an overview of the most common fairness framework, group fairness, and moving to the emerging frameworks of individual and causal fairness [32]. Our discussion includes the mathematical formulation of various fairness criteria, real-world examples, and a case study on mortality prediction using an openly available EHR dataset. We conclude with challenges in existing approaches and prospects for future research. Our review highlights that, in light of the large number of fairness notions, how to operationalize fairness in clinical settings remains an open problem. We therefore aim to provide a concise overview and organization of fairness definitions used within ML in health that can be used as a foundational resource for researchers navigating and contributing to the rich and nuanced fairness literature (see Supplementary Table S1 for a detailed comparison with existing references).
@@
Even with a well-defined and well-intentioned research question, ML models can be unfair due to biases in the data, in the model, and/or in the deployment of the model [33, 34, 35, 36, 25, 37, 38, 39]. Figure 1 provides a visualization of sources of bias that can arise throughout the model development process and Table 1 details common sources of bias together with examples from the health-focused literature.
@@
Figure 1: Sources of bias in developing a machine learning (ML) model. Bias can arise at each stage of the model development process, from data collection to deployment. Each stage influences the next, with bias being potentially perpetuated and compounded throughout the development process. After deployment, an unfair ML can also introduce or reinforce societal bias. This figure is adapted from [40] and [41].
@@
Briefly expanding on Table 1, bias in the data arises from (i) societal and/or (ii) statistical bias [40]. With respect to the former, the data used in health applications most often measures and categorizes people and therefore encodes societal structures, injustices, and stereotypes, such as gender, racial, and age bias [42, 43]. For example, underlying social inequalities in healthcare access can limit the amount of data available in EHRs of certain subgroups, such as Black patients [44]. In terms of statistical bias, the data can also fail to represent the population of interest due the sampling method (selection bias), time of collection (temporal bias), or data quality issues such as missing, mismeasured, or insufficient data (missing data, measurement, and minority bias). In the fairness literature, data with any of these undesirable properties are informally referred to as biased [45, 17]. Observational clinical data (e.g., EHR, medical images) inherently contains various biases and researchers must carefully consider sources of bias in any application [46, 47]. Additionally, choices made during model training and evaluation can further amplify biases in the data or incorporate new bias, including selecting an inappropriate label (label bias), model (algorithmic bias), and/or evaluation metric or data set (evaluation bias). Table 1 details several health examples subject to model bias. Lastly, during deployment, bias in the data and/or model can be reinforced or introduced when users selectively disregard (dismissal bias) or overly trust (automation bias) a model’s outputs, or if models are inaccessible to certain subpopulation (privilege bias). For instance, ML-assisted diagnostic imaging requires that hospitals have the necessary equipment, yet fewer than 20% of rural emergency departments in Canada have in-house computed tomography (CT) scanners [48]. A review of biases in non-health applications can be found in [43] and [34].
@@
Dr. Leo Celi, a health AI expert, emphasizes that “data bias is the waterloo of health AI” and argues that bias must be a primary consideration throughout the ML pipeline - from problem formulation to model deployment [49]. Building on existing AI reporting guidelines, [25] introduced a bias evaluation checklist to enable practitioners to systematically and holistically address bias in clinical predictive models [50, 51]. Bias mitigation strategies and common pitfalls in applying ML in health are further discussed in [52] and [53]. The fairness criteria we introduce in the subsequent sections play an important role in this process. The criteria are typically integrated directly into model training or used as evaluative metrics to identify disparities after the model has been trained [54, 55, 56, 57, 41, 58]. While this review focuses on how to define fairness, strategies for mitigating bias are an equally active and important area of research and have been reviewed from a methodological perspective in [59].
@@
Societal/historical Bias The data reflects long-standing societal disparities encoded within the data over time. Clinical word embeddings trained on large corpora of text, such as clinical notes from healthcare systems, reflect biases about ethnic minorities [60].
Selection Bias The data is not representative of the population of interest. Data in volunteer-based biobanks does not reflect the general population as participation rates are higher among women, older age groups, and individuals with higher socioeconomic status. [61].
Measurement Bias The data contains variables that are collected or measured inaccurately. Gender identity is often recorded in EHRs without provider education and therefore may not accurately represent an individual’s identity [62].
Temporal Bias The data captures a specific time period that may not reflect current or future conditions. Pediatric mental health visits were lower than expected for individuals with lower socioeconomic status during the first year of the COVID-19 pandemic [63].
Minority Bias The data lacks adequate representation from the minority group for the model to accurately learn. Most of the data for genetic studies is from European ancestry populations [64].
Missing Data Bias The data has variables that are incompletely measured. Patients from low-income backgrounds have higher rates of missing medical measurements, which is often due to inequities in access to healthcare [65].
@@
Label Bias An imperfect proxy is selected to train a model instead of the outcome of interest. Healthcare cost was used as a proxy for healthcare need in a commercial algorithm used to identify patients for high-risk care management programs. The algorithm significantly under-identified Black patients for care services as less money is spent on Black patients relative to similarly healthy white patients [16].
Algorithmic Bias Properties of a model and/or its training algorithm create or amplify bias in the data. Differentially private prediction models for mortality performed poorly for minority groups [66].
Evaluation Bias An inappropriate choice of benchmark data or metrics are used for evaluation. Existing public skin disease AI benchmarks do not have images of biopsy-proven malignancy, the gold standard for disease annotation, on dark skin [67, 68].
@@
Automation Bias Model users overly trust model outputs, sometimes even against their own knowledge. A computer-aided diagnosis system that scans mammograms and marks suspicious areas of potential cancer features had lower sensitivity for women aged 40–49 compared to older age groups. A clinician using this system can disproportionately miss cancers for this age group [69].
Dismissal Bias Model users ignore model recommendations, often due to frequent false alerts. If the same computer-aided diagnosis system also has a lower positive predictive value for women aged 40–49, then clinicians may disregard its recommendations, as they are more likely to be false positives [70].
Privilege Bias Models are not available or cannot be deployed in all settings, such as parts of the world where medical technology or sensors are unavailable. Less than 20% of rural emergency departments in Canada have in-house access to computed tomography (CT) scanners [48].
@@
Existing definitions of fairness primarily fall into three categories: group fairness, individual fairness, and causal fairness. Group fairness criteria are commonly used in health and deem a model as fair if its predictions are similarly accurate or calibrated across a predefined set of groups. These groups are most often defined by a protected attribute(s) such as age or race [^2] Sensitive attribute and protected attribute are used interchangeably [71]. Fairness is also evaluated across groups defined by social determinants of health (e.g., income and education) [72]. [34]. Common protected attributes used in the health-focused literature are summarized in Table 2. In contrast, individual fairness is a less commonly used framework and requires that the model provide similar predictions to similar individuals based on a user-defined similarity metric [24, 31]. Lastly, causal fairness criteria utilize causal estimands to quantify unfairness and link observed disparities in model performance to their underlying cause [73, 74]. Causality-based fairness definitions are particularly attractive for health-focused applications as they enable practitioners to interrogate biases [75].
@@
Age [9, 76, 77, 78]
Disability [76]
Marital Status [79]
National Origin [78]
Race [9, 79, 76, 77, 80, 78, 33, 25]
Sex [9, 79, 76, 80, 78]
Socioeconomic Status [9, 79, 77, 78, 33, 25]
@@
Figure 2: Taxonomy of fairness. Fairness criteria primarily fall into three categories: group, individual, and causal fairness. Group fairness criteria are oblivious in the sense that they can be entirely inferred from the distribution of the data. Individual and causal fairness criteria are non-oblivious criteria as they require specification of a similarity metric and a causal model, respectively. Many fairness criteria cannot be simultaneously satisfied and incompatibilities between different notions of fairness are depicted with a red X. This taxonomy is not exhaustive and includes the notions of fairness introduced in this review.
@@
Group fairness criteria require ML models to perform similarly across groups of interest and are the most popular fairness framework in health-focused applications [32]. The criteria primarily fall into three categories: independence, separation, and sufficiency [81, 17]. Table 3 provides mathematical definitions of various criteria and their interpretation in the context of an example. Our discussion focuses on classification as it is most widely used and well-studied [40, 34, 82]. Approaches for continuous outcomes, continuous protected attributes, categorical protected attributes that define more than two groups, and multiple protected attributes (i.e., subgroup fairness) are discussed in Supplementary Sections 2.
@@
Under independence, an ML model is said to be fair if its decisions do not depend on the protected attribute (i.e., $D\perp\!\!\!\perp A$). Statistical (or demographic) parity is a common measure of independence that requires that the model classify individuals into the positive class at the same rate in each group [24, 83]. Conditional statistical parity relaxes this concept by requiring the rate of positive classifications to be the same within more granular groups defined by the protected attribute and other relevant factors. Independence-based metrics are infrequently used in health-focused applications as the prevalence of clinical outcomes often differs across groups (e.g., heart disease in males and females). Enforcing independence may prevent a model from learning a genuine association between the protected attribute and the outcome, which can decrease the overall performance of the model for the entire population [84, 85].
@@
These metrics have been widely used in health-focused applications as incorrect decisions can have severe consequences. The choice of a separation-based metric, however, depends on the specific clinical context. When false negatives have the most severe consequences, equal opportunity may be preferred. For instance, equal opportunity was used in a study of state-of-the-art computer vision models used to detect common diseases from chest X-rays, where a false negative corresponded to incorrectly identifying a patient as not having “no finding” on their X-ray [11]. The models had higher FNRs in several intersectional under-served subpopulations, such as Hispanic female patients, potentially resulting in delayed access to care. There are also situations where balancing both FNR and FPR is more appropriate and equalized odds should be prioritized [54]. For instance, in the context of rapid screening tools for COVID-19 used in emergency departments, a high FNR can lead to insufficient monitoring of COVID-19 while a high FPR can result in unnecessary additional testing [86]. Additional separation-based metrics are detailed in Table 3, including balance for the positive class and balance for the negative class [87].
@@
Sufficiency aims to equalize error rates among individuals with similar decisions [82]. Formally, sufficiency requires that the label does not depend on the protected attribute given the model’s decision ( i.e., $Y\perp\!\!\!\perp A\mid D)$. The decision is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [17]. Common sufficiency-based metrics focus on equalizing the positive predictive value (PPV, known as predictive parity), both the PPV and negative predictive value (NPV, known as conditional use accuracy equality), and calibration (known as well-calibration )[45, 81, 87, 88].
Sufficiency-based metrics are commonly used to evaluate models for disease screening, management, and triage. Conditional use accuracy equality is a particularly common metric, as disparities in PPV and NPV can have distinct impacts on patient care [89, 86]. For example, in COVID-19 triage algorithms, conditional use accuracy seeks to ensure that ICU allocations based on the model’s decisions are determined by medical needs rather than a protected attribute. Balancing the model’s PPV across groups means that patients prioritized for ICU care due to potential severe complications (e.g., high-risk cases) have an equal likelihood of actually experiencing these complications, regardless of protected attributes. Similarly, balancing the NPV across groups ensures that low-risk cases have an equal likelihood of not experiencing complications. Calibration-based metrics are also frequently applied. Well-calibration was used in our introductory example of the commercial risk prediction algorithm used to enroll patients in a high-risk care management program. In this example, healthcare costs were used as a proxy for health needs in training the model. As a result, at equivalent model scores, Black patients were in significantly poorer health than white patients as historically less money had been spent on their healthcare, meaning that they had to be sicker to qualify for the program [16].
@@
Statistical Parity $P(D=1\mid A=a_{0})=P(D=1\mid A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and emergency department (ED) visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females [90].
Conditional Statistical Parity $P(D=1\mid{\bf Z},A=a_{0})=P(D=1\mid{\bf Z},A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and emergency department (ED) visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females after adjusting for age and pre-existing conditions [90].
@@
In predicting asthma exacerbation in children, the rates of false positives (children without exacerbation incorrectly identified as having exacerbation) and false negatives (children with exacerbation incorrectly identified as not having exacerbation) are the same for children from low and high socio-economic classes [91].
Predictive Equality $P(D=1\mid Y=0,A=a_{0})=P(D=1\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting cardiovascular disease (CVD) diagnoses CVD in males and females without CVD at the same rate [92].
Equal Opportunity $P(D=0\mid Y=1,A=a_{0})=P(D=0\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting a suspicious finding from a chest X-rays misses findings in males and females at the same rate [93].
Balance for Positive Class $\mathbb{E}(S\mid Y=1,A=a_{0})=\mathbb{E}(S\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average predicted score of people with lung cancer is equal for males and females [93].
Balance for Negative Class $\mathbb{E}(S\mid Y=0,A=a_{0})=\mathbb{E}(S\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average scores among males and females without lung cancer are equal [93].

### v4

# references.bib

@@ Abstract
There are myriad potential applications of machine learning (ML) for health, including automated disease detection, computer-aided diagnosis, and personalized treatment planning [rajpurkar2022ai]. However, there is substantial evidence that, without appropriate forethought and planning, ML models can introduce or exacerbate health inequities by making less accurate decisions for certain groups or individuals [grote_enabling_2022]. Within medical imaging, state-of-the-art ML models used for disease diagnosis, risk prediction, and triage management are known to underperform within minority groups defined by protected attributes, including sex, race, and ethnicity [rajpurkar_chexnet_2017, ting_development_2017, liu_comparison_2019, lee_machine_2020, seyyed-kalantari_chexclusion_2020, larrazabal_gender_2020, seyyed-kalantari_underdiagnosis_2021, puyol-anton_fairness_2022, bressem_deep_2021, roest_comparative_2022, lin_evaluate_2023]. For example, deep learning models used to detect 14 common diseases from chest X-rays substantially were found to under-diagnose intersectional under-served subgroups, such as Hispanic female patients, potentially resulting in treatment delays if deployed in practice [seyyed-kalantari_underdiagnosis_2021]. Similarly, ML models trained on electronic health record (EHR) and administrative claims data have demonstrated biased performance by underperforming in specific subgroups [li_evaluating_2023, banda_characterizing_2023]. In a landmark study, a widely used commercial risk prediction model for identifying patients with complex health needs had significantly lower accuracy for Black patients than for White patients. This disparity made Black patients less likely than White patients with the same level of illness to be recommended for essential care services [obermeyer_dissecting_2019].
@@
As applications of ML in health become commonplace, it is crucial to recognize, account for, and mitigate such disparities in model performance to ensure health equity. Broadly, an ML model is said to be fair if it does not discriminate against an individual or group [barocas_fairness_2023]. Concepts of fairness have been extensively studied across various disciplines, including social choice theory, game theory, economics, and law [young_equity_1995, roemer_equality_1998, roemer_theories_1998, rawls_theory_1971, rawls_justice_2001, hutchinson201950, dwork_fairness_2012]. Building on these principles, the subfield of fairness within ML provides a framework for evaluating and mitigating bias throughout the model development process.
@@
Although fairness has been an extremely active area of research over the past decade, implementing fairness within ML for health is relatively nascent [wang_bias_2022, feng_fair_2022, liu_translational_2023]. A systematic review of articles utilizing ML for EHR-based phenotyping found that only 5% of studies assessed fairness [yang_machine_2023]. Similarly, a review of EHR-based prediction models found that most studies investigate the overall performance of ML models, but do not interrogate potential biases [chen2024unmasking]. Beyond EHR applications, several scoping reviews of clinical ML models found that the adoption of fairness remains inconsistent, partly due to a knowledge gap between ML and clinical researchers [liu_translational_2023, plecko_causal_2022, anderson2024algorithmic]. Moreover, over two dozen definitions of fairness have been proposed, most of which originate outside health-focused literature, making it a particularly challenging domain to navigate.
@@
To review this fundamental topic, we examine key notions of fairness and their use in ML in health [^1] This review focuses on quantitative measures of fairness. Alternative approaches are discussed in the outlook section.. We first introduce biases that can emerge throughout the model development process using examples from the literature to illustrate why ML models can be unfair. We then review what it means for a model to be fair, beginning with an overview of the most common fairness framework, group fairness, and moving to the emerging frameworks of individual and causal fairness [awasthi2020beyond]. Our discussion includes the mathematical formulation of various fairness criteria, real-world examples, and a case study on mortality prediction using an openly available EHR dataset. We conclude with challenges in existing approaches and prospects for future research. Our review highlights that, in light of the large number of fairness notions, how to operationalize fairness in clinical settings remains an open problem. We therefore aim to provide a concise overview and organization of fairness definitions used within ML in health that can be used as a foundational resource for researchers navigating and contributing to the rich and nuanced fairness literature (see Supplementary Table S1 for a detailed comparison with existing references).
@@
Even with a well-defined and well-intentioned research question, ML models can be unfair due to biases in the data, in the model, and/or in the deployment of the model [rajkomar_ensuring_2018, mehrabi_survey_2021, fazelpour_algorithmic_2021, mhasawade2021machine, wang_bias_2022, xu_algorithmic_2022, drukker_toward_2023, ferrara_fairness_2024]. Figure 1 provides a visualization of sources of bias that can arise throughout model development and Table 1 details common biases using examples from the literature.
@@
Figure 1: Sources of bias. Bias can arise at each stage of the model development process, from data collection to deployment. Each stage influences the next, with bias being potentially perpetuated and compounded throughout the development process. After deployment, an unfair model can also introduce or reinforce societal bias. This figure is adapted from [mitchell_algorithmic_2021] and [mbakwe2023fairness].
@@
Briefly expanding on Table 1, bias in the data arises from (i) societal and/or (ii) statistical bias [mitchell_algorithmic_2021]. With respect to the former, the data used in health applications most often measures and categorizes people and therefore encodes societal structures, injustices, and stereotypes, such as gender, racial, and age bias [iliadis2016critical, suresh2021framework]. For example, underlying social inequalities in healthcare access can limit the amount of data available in EHRs of certain subgroups, such as Black patients [getzen_mining_2023]. In terms of statistical bias, the data can also fail to represent the population of interest due the sampling method (selection bias), time of collection (temporal bias), or data quality issues such as missing, mismeasured, or insufficient data (missing data, measurement, and minority bias). In the fairness literature, data with any of these undesirable properties are informally referred to as biased [chouldechova_fair_2017, barocas_fairness_2023]. Observational clinical data (e.g., EHR, medical images) inherently contains various biases and researchers must carefully consider sources of bias in any application [genevieve_structural_2020, mccoy_ensuring_2020]. Additionally, choices made during model training and evaluation can further amplify biases in the data or incorporate new bias, including selecting an inappropriate label (label bias), model (algorithmic bias), and/or evaluation metric or data set (evaluation bias). Table 1 details several health examples subject to model bias. Lastly, during deployment, bias in the data and/or model can be reinforced or introduced when users selectively disregard (dismissal bias) or overly trust (automation bias) a model’s outputs, or if models are inaccessible to certain subpopulation (privilege bias). For instance, ML-assisted diagnostic imaging requires that hospitals have the necessary equipment, yet fewer than 20% of rural emergency departments in Canada have in-house computed tomography (CT) scanners to utilize these advances [davidson_landscape_2024]. A review of biases outside of health-focused applications can be found in [suresh2021framework] and [mehrabi_survey_2021].
@@
Dr. Leo Celi, a health AI expert, emphasizes that “data bias is the waterloo of health AI.” Bias must be a primary consideration throughout the ML pipeline - from problem formulation to model deployment [tedx_talks_data_2023]. Building on existing AI reporting guidelines, [wang_bias_2022] introduced a bias evaluation checklist to enable practitioners to systematically and holistically address bias in clinical predictive models [wolff2019probast, liu2020reporting]. Bias mitigation strategies and common pitfalls in applying ML in health are further discussed in [gichoya2023ai] and [panch_artificial_2019]. The fairness criteria we introduce in the subsequent sections play an important role in this process. The criteria are typically integrated directly into model training or used as evaluative metrics to identify disparities after the model has been trained [hardt_equality_2016, menon_cost_2018, pessach_review_2022, brahmbhatt2023towards, mbakwe2023fairness, goethals2024beyond]. While this review focuses on how to define fairness, strategies for mitigating bias are an equally important area of research and have been reviewed in [hort2023bias].
@@
Societal (or historical) Bias The data reflects long-standing societal disparities encoded within the data over time. Clinical word embeddings trained on large corpora of text, such as clinical notes from healthcare systems, reflect biases about ethnic minorities [zhang2020hurtful].
Selection Bias The data is not representative of the population of interest. Data in volunteer-based biobanks does not reflect the general population as participation rates are higher among women, older age groups, and individuals with higher socioeconomic status. [fry_comparison_2017].
Measurement Bias The data contains variables that are collected or measured inaccurately. Gender identity is often recorded in EHRs without provider education and therefore may not accurately represent an individual’s identity [kronk_transgender_2022].
Temporal Bias The data captures a specific time period that may not reflect current or future conditions. Pediatric mental health visits were lower than expected for individuals with lower socioeconomic status during the first year of the COVID-19 pandemic [toulany_sociodemographic_2023].
Minority Bias The data lacks adequate representation from the minority group for the model to accurately learn. Most of the data for genetic studies is from European ancestry populations [martin_clinical_2019].
Missing Data Bias The data has variables that are incompletely measured. Patients from low-income backgrounds have higher rates of missing medical measurements, which is often due to inequities in access to healthcare [Chapman_Exploring_2022].
@@
Label Bias An imperfect proxy is selected to train a model instead of the outcome of interest. Healthcare cost was used as a proxy for healthcare need in a commercial algorithm used to identify patients for high-risk care management programs. The algorithm significantly under-identified Black patients for care services as less money is spent on Black patients relative to similarly healthy white patients [obermeyer_dissecting_2019].
Algorithmic Bias Properties of a model and/or its training algorithm create or amplify bias in the data. Differentially private prediction models for mortality performed poorly for minority groups [suriyakumar2021chasing].
Evaluation Bias An inappropriate choice of benchmark data or metrics are used for evaluation. Existing public skin disease AI benchmarks do not have images of biopsy-proven malignancy, the gold standard for disease annotation, on dark skin [daneshjou_lack_2021, daneshjou_disparities_2022].
@@
Automation Bias Model users overly trust model outputs, sometimes even against their own knowledge. A computer-aided diagnosis system that scans mammograms and marks suspicious areas of potential cancer features had lower sensitivity for women aged 40–49 compared to older age groups. A clinician using this system can disproportionately miss cancers for this age group [lehman_diagnostic_2015].
Dismissal Bias Model users ignore model recommendations, often due to frequent false alerts. If the same computer-aided diagnosis system also has a lower positive predictive value for women aged 40–49, then clinicians may disregard its recommendations, as they are more likely to be false positives [ancker2017effects].
Privilege Bias Models are not available or cannot be deployed in all settings, such as parts of the world where medical technology is unavailable. Less than 20% of rural emergency departments in Canada have in-house access to computed tomography (CT) scanners required for ML-assisted diagnostic imaging [davidson_landscape_2024].
@@
Existing definitions of fairness primarily fall into three categories: group fairness, individual fairness, and causal fairness. Group fairness criteria are commonly used in health and deem a model as fair if its predictions are similarly accurate or calibrated across a predefined set of groups. These groups are most often defined by a protected attribute(s) such as age or race [^2] Sensitive attribute and protected attribute are used interchangeably [caton_fairness_2024]. Fairness is also evaluated across groups defined by social determinants of health (e.g., income and education) [mccradden2020ethical]. [mehrabi_survey_2021]. Common protected attributes used in the health-focused literature are summarized in Table 2. In contrast, individual fairness is a less commonly used framework and requires that the model provide similar predictions to similar individuals based on a user-defined similarity metric [dwork_fairness_2012, anderson2024algorithmic]. Lastly, causal fairness criteria utilize causal estimands to quantify unfairness and link observed disparities in model performance to their underlying cause [plecko2022causal, makhlouf_when_2024]. Causality-based fairness definitions are particularly attractive for health-focused applications as they enable practitioners to interrogate biases [zhang_causal_2024].
@@
Age [seyyed-kalantari_chexclusion_2020, garriga_machine_2022, schuch_fairness_2023, yfantidou_accuracy_2023]
Disability [garriga_machine_2022]
Marital Status [wang_comparison_2022]
National Origin [yfantidou_accuracy_2023]
Race [seyyed-kalantari_chexclusion_2020, wang_comparison_2022, garriga_machine_2022, schuch_fairness_2023, yuan_algorithmic_2023, yfantidou_accuracy_2023, rajkomar_ensuring_2018, wang_bias_2022]
Sex [seyyed-kalantari_chexclusion_2020, wang_comparison_2022, garriga_machine_2022, yuan_algorithmic_2023, yfantidou_accuracy_2023]
Socioeconomic Status [seyyed-kalantari_chexclusion_2020, wang_comparison_2022, schuch_fairness_2023, yfantidou_accuracy_2023, rajkomar_ensuring_2018, wang_bias_2022]
@@
Figure 2: Taxonomy of fairness. Fairness criteria primarily fall into three categories: group, individual, and causal fairness. Group fairness criteria are oblivious in the sense that they can be entirely inferred from the distribution of the data. Individual and causal fairness criteria are non-oblivious criteria as they require specification of similarity metrics and a causal model, respectively. Many fairness criteria cannot be simultaneously satisfied and incompatibilities between different notions of fairness are depicted with a red X. This taxonomy is not exhaustive and includes the notions of fairness introduced in this review.
@@
Group fairness criteria require ML models to perform similarly across groups of interest and are the most popular fairness framework in health-focused applications [awasthi2020beyond]. The criteria primarily fall into three categories: independence, separation, and sufficiency [berk_fairness_2018, barocas_fairness_2023]. Table 3 provides mathematical definitions of various criteria and their interpretation in the context of an example. Our discussion focuses on classification as it is most widely used and well-studied [mitchell_algorithmic_2021, mehrabi_survey_2021, castelnovo_clarification_2022]. Approaches for continuous outcomes, continuous protected attributes, categorical protected attributes that define more than two groups, and multiple protected attributes (i.e., subgroup fairness) are discussed in Supplementary Section 2.
@@
Under independence, an ML model is said to be fair if its decisions do not depend on the protected attribute (i.e., $D\perp\!\!\!\perp A$). Statistical (or demographic) parity is a common measure of independence that requires that the model classify individuals into the positive class at the same rate in each group [dwork_fairness_2012, corbett-davies_algorithmic_2017]. Conditional statistical parity relaxes this concept by requiring the rate of positive classifications to be the same within more granular groups defined by the protected attribute and other relevant factors. Independence-based metrics are infrequently used in health-focused applications as the prevalence of clinical outcomes often differs across groups (e.g., heart disease in males and females). Enforcing independence may also prevent a model from learning a genuine association between the protected attribute and the outcome, which can decrease the overall performance of the model for the entire population [pfohl2019creating, lei2024inductive].
@@
Separation-based metrics have been widely used in health-focused applications, although the specific choice of metric depends on the clinical context. When false negatives have the most severe consequences, equal opportunity may be preferred. For instance, equal opportunity was used in a study of state-of-the-art computer vision models used to detect common diseases from chest X-rays, where a false negative corresponded to incorrectly identifying a patient as not having “no finding” on their X-ray [seyyed-kalantari_underdiagnosis_2021]. The models had higher FNRs in several intersectional under-served subpopulations, such as Hispanic female patients, potentially resulting in delayed access to care. There are also situations where balancing both the FNR and FPR is more appropriate and equalized odds should be prioritized [hardt_equality_2016]. For instance, in the context of rapid screening tools for COVID-19 used in emergency departments, a high FNR can lead to insufficient monitoring of COVID-19 while a high FPR can result in unnecessary additional testing in certain groups [yang2023algorithmic]. Additional separation-based metrics are detailed in Table 3, including balance for the positive class and balance for the negative class [kleinberg2017inherent].
@@
Sufficiency aims to equalize error rates among individuals with similar decisions [castelnovo_clarification_2022]. Formally, sufficiency requires that the label does not depend on the protected attribute given the model’s decision ( i.e., $Y\perp\!\!\!\perp A\mid D)$. The decision is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [barocas_fairness_2023]. Common sufficiency-based metrics focus on equalizing the positive predictive value (PPV, known as predictive parity), both the PPV and negative predictive value (NPV, known as conditional use accuracy equality), and calibration (known as well-calibration )[chouldechova_fair_2017, berk_fairness_2018, kleinberg2017inherent, guo_calibration_2017].
Sufficiency-based metrics are often used to evaluate models for disease screening, management, and triage. Conditional use accuracy equality is a particularly common metric, as disparities in PPV and NPV can have distinct impacts on patient care [brandao-de-resende_machine_2023, yang2023algorithmic]. For example, in COVID-19 triage algorithms, conditional use accuracy seeks to ensure that ICU allocations based on the model’s decisions are determined by medical needs rather than a protected attribute. Balancing the model’s PPV across groups means that patients prioritized for ICU care due to potential severe complications (e.g., high-risk cases) have an equal likelihood of actually experiencing these complications, regardless of protected attributes. Similarly, balancing the NPV across groups ensures that low-risk cases have an equal likelihood of not experiencing complications. Calibration-based metrics are also frequently applied. Well-calibration was evaluated in our introductory example of the commercial risk prediction algorithm used to enroll patients in a high-risk care management program. In this example, healthcare costs were used as a proxy for health needs in training the model. As a result, at equivalent model scores, Black patients were in significantly poorer health than white patients as historically less money had been spent on their healthcare, meaning that they had to be sicker to qualify for the program [obermeyer_dissecting_2019].
@@
Statistical Parity $P(D=1\mid A=a_{0})=P(D=1\mid A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and emergency department (ED) visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females [davoudi_fairness_2024].
Conditional Statistical Parity $P(D=1\mid{\bf Z},A=a_{0})=P(D=1\mid{\bf Z},A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and ED visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females after adjusting for age and pre-existing conditions [davoudi_fairness_2024].
@@
In predicting asthma exacerbation in children, the rates of false positives (children without exacerbation incorrectly identified as having exacerbation) and false negatives (children with exacerbation incorrectly identified as not having exacerbation) are the same for children from low and high socio-economic classes [juhn2022assessing].
Predictive Equality $P(D=1\mid Y=0,A=a_{0})=P(D=1\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting cardiovascular disease (CVD) diagnoses CVD in males and females without CVD at the same rate [Vargae002395].
Equal Opportunity $P(D=0\mid Y=1,A=a_{0})=P(D=0\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting a suspicious finding from a chest X-rays misses findings in males and females at the same rate [ueda_artificial_2021].
Balance for Positive Class $\mathbb{E}(S\mid Y=1,A=a_{0})=\mathbb{E}(S\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average score of people with lung cancer is equal for males and females [ueda_artificial_2021].
Balance for Negative Class $\mathbb{E}(S\mid Y=0,A=a_{0})=\mathbb{E}(S\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average score among males and females without lung cancer are equal [ueda_artificial_2021].